OpenStudy (nuccioreggie):

Bella, Lydia, Camila, and Trinity together stacked a total of 60 pizzas in a pan. Bella stacked 0.25 of the pizzas, Lydia stacked 12 of the 60 pizzas, Camila stacked 35% of the pizzas, and Trinity stacked the remaining pizzas. Who stacked the most pizzas? Bella Lydia Camila Trinity

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (anonymous):

ok, first you need to find out how you're measuring this. percentages, decimals, or numbers. Then add up all that and find the remaining for Trinity. Then you choose who has the greatest number/decimal/percentage HOpe this helps

OpenStudy (nuccioreggie):

Ok i thought the answer was b

OpenStudy (anonymous):

hang on, let me c...

OpenStudy (nuccioreggie):

Ok

OpenStudy (anonymous):

ok, so here's what I got: Bella - 1/4 Lydia - 1/5 Camila - 7/20 Trinity - 1/5 Who has the most?

OpenStudy (anonymous):

if this helps: Bella - 5/20 Lydia - 4/20 Oh, and you multiply each by 3/1 because the total number of pizzas is 60

OpenStudy (nuccioreggie):

So it was bella???

OpenStudy (nuccioreggie):

@Owl_eyes2

OpenStudy (michele_laino):

hint: 0.25 of 60, means 60/4, so Bella stacked: 60/4 =...? pizzas

OpenStudy (michele_laino):

35% of 60 means 60 * 35/100=...? so Camila stacked 60*35/100=...? pizzas

OpenStudy (nuccioreggie):

15

OpenStudy (michele_laino):

ok! and what is: 60*35/100=...?

OpenStudy (nuccioreggie):

21

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

reassuming: Bella stacked 15 pizzas Camila stacked 21 pizzas ydia stacked 12 pizzas and trinity stacked 60-(15+21+12)=...?

OpenStudy (michele_laino):

oops..Lydia*

OpenStudy (nuccioreggie):

78

OpenStudy (michele_laino):

what is: 15+21+12=...?

OpenStudy (nuccioreggie):

48

OpenStudy (michele_laino):

ok! and what is: 60-48=...?

OpenStudy (michele_laino):

@nuccioreggie

OpenStudy (nuccioreggie):

12

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

Hellooooooooo are u there brooo hellooo

OpenStudy (michele_laino):

that's right! So what is the right option?

OpenStudy (michele_laino):

@nuccioreggie